OpenStudy (animalover02):

I WILL METAL!!!! NEED THE GRAPH FOR THIS! please which graph is 3y-5x is less than or equal to -6?

OpenStudy (animalover02):

i will metal...:)

563blackghost (563blackghost):

Convert to slope-intercept form \(\huge\bf{3y - 5x \le -6}\) First add 5x to each side. \(\huge\bf{3y \le 5x - 6}\) Now divide by 3 to isolate y. \(\huge\bf{y=\frac{5}{3}x-2}\)

563blackghost (563blackghost):

Ooops not equal I meant \(\huge\bf{y \le \frac{5}{3}x - 2}\)

OpenStudy (animalover02):

can you do it like you did the first question earlier got me the answer by graph please. this test is going to shut off in 2 min. and im not very fast.

OpenStudy (animalover02):

@Will.H

OpenStudy (animalover02):

@563blackghost

OpenStudy (will.h):

using the data given by ghosty the inequality will have a complete sign (not dached) with a slope of 5/3 and y intercept of (0,-2) and shaded to the right bottom corner as shown in the attachment

OpenStudy (animalover02):

and the line has to be dotted or solid? @Will.H @563blackghost

OpenStudy (will.h):

i mentioned above Not dashed.. means solid

OpenStudy (animalover02):

ok thanks

OpenStudy (will.h):

because of the equal sign

OpenStudy (will.h):

you welcome